NASA Space Station On-Orbit Status 29 January 2018 - Focus Shifts to Russian Spacewalk Friday

Cosmonauts Anton Shkaplerov (left) and Alexander Misurkin work inside the U.S. Destiny laboratory module participating in eye exams using a fundoscope. Credit: NASA.

Expedition 54 is now focusing on Friday's spacewalk to install and remove gear on the Russian side of the International Space Station. This comes after today's spacewalk to work on the Canadarm2 robotic arm was postponed to mid-February.

Cosmonauts Alexander Misurkin and Anton Shkaplerov will put on their Orlan spacesuits Friday and exit the Pirs airlock around 10:30 a.m. EST. The duo will be spacewalking for about six and a half hours to replace an electronics box for a communications antenna on the aft end of the Zvezda service module. The old box will be jettisoned into space to eventually burn up in the Earth's atmosphere.

If time allows, the Russian spacewalkers may be able work a few more tasks. A pair of exposed experiments, Test and Biorisk, are due to be retrieved and brought back inside the station. The cosmonauts may also photograph the back of Zvezda, reposition a foot restraint and jettison old experiment gear.

Over the weekend space station managers postponed today's spacewalk for robotics maintenance. Astronauts Mark Vande Hei and Norishige Kanai were expected to swap a Latching End Effector (LEE) on the tip of the Canadarm2 today installed last week on a spacewalk; however, Canadian Space Agency engineers uploaded a diagnostics software to identify the primary communications string anomaly and test the software upgrade on the spare LEE, eliminating today's swap work. The final software patch will be uploaded in early February.

The duo is now planning for a mid-February spacewalk to bring an end effector inside the station removed from the arm during a spacewalk last October, and install the end effector removed last Tuesday on the mobile base system rail car on the station's truss.

On-Orbit Status Report

USOS ExtraVehicular Activity (EVA) 48 status: Over the weekend, ground teams were able to identify a root cause and a fix for the Space Station Remote Manipulator System (SSRMS) Latching End Effector (LEE) anomaly seen after USOS EVA 47. Because of this, the EVA to return the LEE to the configuration prior to EVA 47 has been deferred. EVA 48 is now targeted for mid-February, and the tasks for that EVA are currently being replanned.

SSRMS LEE status: During USOS EVA 47, LEE serial number 201 was removed from the SSRMS and replaced with the spare LEE, serial number 204. After the installation, ground teams were unable to bring LEE 204 to an operational state on primary power. Redundant power operated as expected. Due to this anomaly, plans were put in place to swap the LEEs back and place LEE 201 on the SSRMS as it was before. Over the weekend, ground teams were able to install diagnostic software that identified the root cause as a faulty initial force reading. Ground and on orbit testing confirmed the root cause, and also confirmed the ability to operate the LEE on the prime power string if the LEE was already grappled to a fixture. A software update is being developed to resolve the issue, and is targeted for uplink in two weeks, at which point it is expected that the LEE currently installed on the SSRMS (LEE 204) will be fully operational.

Electro-static Levitation Furnace (ELF) sample transfer: Today the crew moved samples within ELF to prepare for upcoming ground commanded operations. The Electrostatic Levitation Furnace (ELF) is an experimental facility designed to levitate/melt/solidify materials by containerless processing techniques using the Electrostatic Levitation method. With this facility, thermophysical properties of high temperature melts can be measured, and solidification from deeply undercooled melts can be achieved. The ELF is located in the JEM Multipurpose Small Payload Rack (MSPR) in Kibo.

At Home in Space: The crew completed an At Home in Space questionnaire this morning. This Canadian Space Agency experiment assesses culture, values, and psychosocial adaptation of astronauts to a space environment shared by multinational crews on long-duration missions. It is hypothesized that astronauts develop a shared space culture that is an adaptive strategy for handling cultural differences and they deal with the isolated confined environment of the spacecraft by creating a home in space. At Home in Space uses a questionnaire to investigate individual and culturally related differences, family functioning, values, coping with stress, and post-experience growth.

EPO: Today the crew recorded several demonstration videos to explain different aspects of life in space to schoolchildren on the ground. The first video completed today demonstrates the way that crewmembers sleep in space, while the second demonstrates the challenges of and solutions used for exercising the human body while on orbit.

Lighting Effects: The crew conducted a Visual Performance Test by stowing the hardware in their crew quarters, setting the light to the correct mode, turning all other light sources in the crew quarters off, and performing a Color Discrimination Test. The Lighting Effects investigation studies the impact of the change from fluorescent light bulbs to solid-state light-emitting diodes (LEDs) with adjustable intensity and color and aims to determine if the new lights can improve crew circadian rhythms, sleep, and cognitive performance.

Permanent Multipurpose Module (PMM) rack moves: In order to increase efficiency in the PMM, today the crew moved several stowage racks inside the PMM. The crew moved a Zero-G Stowage Rack (ZSR) and Resupply Stowage Platform (RSP), allowing easier accessibility to the stowage in these racks.

Temperature and Humidity Control (THC) High-Efficiency Particulate Air (HEPA) filter R&R: Today the crew removed and replaced HEPA filters in the JEM. This is regularly schedule maintenance to keep airflow high and particulate contamination low.

Today's Planned Activities
All activities are complete unless otherwise noted.

Big Picture Word Review for PMM Rack Swap Activities
Electrostatic Levitation Furnace(ELF) Sample Transfer On-board Training
JEM DMS1 Rack (JPM1O5) Cable Reconfiguration
On MCC Go Regeneration of БМП Ф1 Micropurification Cartridge (init)
JEM Stowage Consolidation Part 1
Electrostatic Levitation Furnace(ELF) sample Cartridge Retrieval
Electrostatic Levitation Furnace(ELF) sample Holder Removal
Electrostatic Levitation Furnace(ELF) Sample Refill
Water Recovery System Waste Water Tank Drain Init
SPHERES Battery Setup
Water Recovery System (WRS) Waste Water (WW) Tank Sample and Drain Termination
Glacier Desiccant Swap
ПФ1, ПФ2 Dust Filter Cleaning and DC1 В1, В2 Fan Grille Cleaning
Station Support Computer 23 Hard Disk Drive Swap
Electrostatic Levitation Furnace(ELF) Sample Holder Installation
Vacuum cleaning of ВД1 and ВД2 air ducts in DC1
Electrostatic Levitation Furnace(ELF) sample Cartridge Installation
At Home In Space Questionnaire
Robotic Workstation (RWS) Teardown
В3 fan grill cleaning in DC1
File downlink from Soyuz 734 Samsung PC per ODF DESCENT
Extravehicular Activity PGT Battery Removal
Extravehicular Activity (EVA) Airlock Restow
Structures and Mechananisms Permanent Multipurpose Module (PMM) Rack Reconfiguration Part 1
CWC-I to EDV transfer init
Record of EPO Paxi script
SPHERES Battery Swap
JPM THC A High Efficiency Particulate Air (HEPA) Filter R&R
СОЖ maintenance
JPM THC B High Efficiency Particulate Air (HEPA) Filter R&R
CWC-I to EDV transfer term
Crew Medical Officer (CMO) On Board Training (OBT)
БД-2 harness carrier adjustment and equipment restart
Lighting Effects Visual Performance Tests Light Setting - Subject
Water Resource Management (WRM) Condensate Sample Init
JPM THC A&B High Efficiency Particulate Air (HEPA) Filter Closeout
Exercise Data Downlink via OCA
Structures and Mechanisms Permanent Multipurpose Module (PMM) Rack Reconfiguration Part 2
Water Resource Management (WRM) Condensate Sample Terminate
On MCC Go Regeneration of БМП Ф1 Absorption Cartridge (term)

Completed Task List Activities
Manufacturing Device ABS Print Removal, Clean and Stow
WHC KTO Replace
Ziplock Restow

Ground Activities
All activities are complete unless otherwise noted.
Standard commanding.

Three-Day Look Ahead:
Tuesday, 01/30: RS EVA assist, RR habitat restock, CWSA DM R&R
Wednesday, 01/31: Eye exams
Thursday, 02/01: ACE oil dispensing, eye exams

QUICK ISS Status - Environmental Control Group:
Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - Off
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Operate
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Standby
Major Constituent Analyzer (MCA) Lab - Idle
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Process
Trace Contaminant Control System (TCCS) Lab - Full Up
Trace Contaminant Control System (TCCS) Node 3 - Off

Please follow SpaceRef on Twitter and Like us on Facebook.